# DBAW-PIKAN: Dynamic Balance Adaptive Weight Kolmogorov-Arnold Neural Network for Solving Partial Differential Equations

### v1

@@ 2.3 Dynamic Adaptive Loss Weighting Function
In Physics-Informed Neural Networks (PINNs), although they effectively solve many simple problems, they often encounter difficulties when dealing with more complex ones. Traditional PINN methods use fixed weights to balance these loss terms, but their performance is highly dependent on the choice of weights and requires tedious manual tuning. Therefore, selecting appropriate loss weights to balance the various components of the loss function can effectively accelerate the convergence speed and improve the accuracy of PINNs[mcclenny2023selfadaptive][xiang2022selfadaptive]. Inspired by the uncertainty weighting method in multi-task learning[kendall2018multitask], we can treat each loss term—which typically includes the Partial Differential Equation (PDE) residual, Boundary Conditions (BC), and Initial Conditions (IC)—as a separate task. Thus, the process of solving a PDE can be viewed as a multi-task learning problem. By using homoscedastic uncertainty as the basis for loss weighting in a multi-task learning problem, we note that homoscedastic uncertainty is an impromptu uncertainty that does not depend on the input data. It is not a model output but a quantity that remains constant across all input data and varies between different tasks. Therefore, it can be described as task-dependent uncertainty. The implementation process is as follows. We assume that the PDE solution follows a Gaussian distribution, and its likelihood function can be expressed as:
$p(u|\hat{u},\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\|u-\hat{u}\|^{2}}{2\sigma^{2}}\right)$ (14)
@@
$p(u|\hat{u},\sigma_{r}^{2},\sigma_{b}^{2},\sigma_{i}^{2})=\prod_{j\in\{r,b,i\}}\frac{1}{\sqrt{2\pi\sigma_{j}^{2}}}\exp\left(-\frac{\mathcal{L}_{j}}{2\sigma_{j}^{2}}\right)$ (15)
@@
$\begin{split}&-\log p(u|\hat{u},\sigma_{r}^{2},\sigma_{bc}^{2},\sigma_{ic}^{2})\\ &\quad=-\sum_{j\in\{r,bc,ic\}}\log\biggl[\frac{1}{\sqrt{2\pi\sigma_{j}^{2}}}\exp\biggl(-\frac{\|u-\hat{u}\|^{2}}{2\sigma_{j}^{2}}\biggr)\biggr]\\ &\quad=\sum_{j\in\{r,bc,ic\}}\biggl(\frac{1}{2\sigma_{j}^{2}}\mathcal{L}_{j}+\frac{1}{2}\log\sigma_{j}^{2}+\frac{1}{2}\log(2\pi)\biggr)\end{split}$ (16)
@@
$\mathcal{L}_{\text{base}}(W,\sigma_{r}^{2},\sigma_{b}^{2},\sigma_{i}^{2})=\sum_{j\in\{r,b,i\}}\left(\frac{1}{2\sigma_{j}^{2}}\mathcal{L}_{j}(W)+\frac{1}{2}\log\sigma_{j}^{2}\right)$ (17)
The first term in eq17 represents the weighted loss function for each task, while the second term is a regularization term that penalizes tasks with high uncertainty. Minimizing $\mathcal{L}_{\text{base}}$ in eq17 is equivalent to maximizing the joint probability distribution of all losses, which facilitates the adaptive allocation of appropriate weight coefficients to the various components of the loss, ensuring a balance among them. The parameter $\sigma_{j}^{2}$ represents the uncertainty associated with each task. If a task is easy to learn and exhibits low uncertainty, it will be assigned a higher weight; conversely, a task with high uncertainty will receive a lower weight. In summary, the uncertainty-based loss function is:
@@
$\quad+\frac{1}{2}\log\sigma_{r}^{2}+\frac{1}{2}\log\sigma_{bc}^{2}+\frac{1}{2}\log\sigma_{ic}^{2}$
However, in multi-task PDE solving, different loss terms (PDE residual, boundary conditions, initial conditions) often have different orders of magnitude and convergence characteristics. In the early stages of training, certain loss terms may dominate the gradient updates. If the gradient magnitude of some loss terms is too large, it can lead to training instability and weight imbalance, causing the weights of some tasks to grow excessively, thereby suppressing the learning of other tasks and causing the model to prematurely fall into local optima for certain tasks. Therefore, inspired by [niu2025improved], this paper introduces a dynamic upper bound weight constraint method. Its core mathematical expression is:
$\gamma(t)=\gamma_{\text{max}}\cdot\exp(-\alpha t)+\gamma_{\text{min}}$ (19)
@@
$\lambda_{j}=\frac{1}{\sigma_{j}^{2}+\frac{1}{\gamma(t)}+\epsilon}$ (20)
where $\epsilon=10e^{-12}$ is a small constant added for numerical stability. When $\gamma(t)$ is large, $\frac{1}{\gamma(t)}\rightarrow 0$, and the weight is mainly controlled by $\sigma_{j}^{2}$; when $\gamma(t)$ is small, the term $\frac{1}{\gamma(t)}$ becomes significant, constraining the range of weight changes. Finally, we obtain the extended constrained adaptive loss function:
$\mathcal{L}_{\text{adaptive}}=\sum_{j\in\{r,b,i\}}\left(\lambda_{j}\mathcal{L}_{j}+\log\left(\sigma_{j}^{2}+\frac{1}{\gamma(t)}\right)\right)$ (21)
@@ 2.4 DBAW-PIKAN
To simultaneously address the limitations in network expressivity and the imbalance in multi-objective optimization within the PINN framework, this study proposes an integrated DBAW-PIKAN framework. This framework innovatively consolidates the high-expressivity PIKAN architecture described in Section 2.2 with the DBAW self-adaptive weighting strategy introduced in Section 2.3.
This integration is not a mere superposition of two independent improvements but is fundamentally rooted in the synergy between them.On one hand, the PIKAN architecture, as constructed in Section 2.2, leverages the superior function approximation capabilities of Kolmogorov-Arnold Networks to provide the potential for accurate PDE solutions, effectively mitigating the spectral bias inherent in traditional MLPs [wang2025kolmogorov]. However, the potential of a high-expressivity architecture cannot be fully realized if the training process is unstable. As an implementation of PINNs, PIKAN still inherits the core challenge of imbalanced training dynamics caused by multi-objective loss functions.On the other hand, the DBAW strategy proposed in Section 2.3 provides stability to the optimization process by dynamically balancing the contributions of various loss components. This effectively prevents certain loss terms from dominating the optimization and suppressing the learning of other physical constraints. However, experimental results (see Section 3.2) indicate that the DBAW strategy is not a universal ”plug-and-play” module. When DBAW is combined with a traditional MLP architecture (i.e., the DBAW-PINN model), it exhibits suboptimal performance in solving the nonlinear Burgers equation. This suggests that the efficacy of the DBAW strategy is highly dependent on a compatible and sufficiently expressive underlying network architecture.
@@
Input: $\gamma_{\max}$, $\gamma_{\min}$, $\alpha$, $N_{\text{int}}$, $N_{\text{bc}}$, $T$, epoch
@@
1:Initialize KAN model $f_{\theta}$
2:Initialize adaptive weight parameters $\log\sigma_{r}\leftarrow 0$, $\log\sigma_{u}\leftarrow 0$
3:Initialize optimizers Adamθ and Adamσ with learning rates
4:for $t=1,\ldots,T$ do
5:  Sample interior points $\{(x_{i},y_{i})\}_{i=1}^{N_{\text{int}}}$ and boundary points $\{(x_{j}^{b},y_{j}^{b})\}_{j=1}^{N_{\text{bc}}}$
6:  Compute dynamic upper bound: $\gamma(t)=\gamma_{\max}\cdot\exp(-\alpha t)+\gamma_{\min}$
7:  Compute PDE residual loss $\mathcal{L}_{\text{PDE}}$ and boundary loss $\mathcal{L}_{\text{BC}}$
8:  Calculate adaptive weights:
9:   $\sigma_{r}=\exp(\log\sigma_{r})$, $\sigma_{u}=\exp(\log\sigma_{u})$
10:   $\lambda_{r}=\min\left\{\frac{1}{\sigma_{r}^{2}+\frac{1}{\gamma(t)}},\gamma(t)\right\}$
11:   $\lambda_{u}=\min\left\{\frac{1}{\sigma_{u}^{2}+\frac{1}{\gamma(t)}},\gamma(t)\right\}$
12:  Compute total loss with regularization and smoothness penalty
13:  $\theta_{t+1}=\text{Adam}_{\theta}(\mathcal{L}_{\text{total}},\theta_{t})$
14:  $\log\sigma_{r,t+1},\log\sigma_{u,t+1}=\text{Adam}_{\sigma}(\mathcal{L}_{\text{total}},\log\sigma_{r,t},\log\sigma_{u,t})$
15:end for
16:return $\theta^{*}$, $\log\sigma_{r}^{*}$, $\log\sigma_{u}^{*}$
@@ 3 Experiments
To quantify the prediction error of the models, this study employs the relative $L_{2}$ error as the performance metric, which is defined as:
$L_{2}=\frac{\sqrt{\sum_{i=1}^{N}\left|\hat{u}\left(x_{i},t_{i}\right)-u\left(x_{i},t_{i}\right)\right|^{2}}}{\sqrt{\sum_{i=1}^{N}\left|u\left(x_{i},t_{i}\right)\right|^{2}}}$ (22)
@@
The objective of this section is to evaluate the performance of the proposed DBAW self-adaptive loss strategy and the enhanced PIKAN framework through a series of numerical experiments. The experimental design encompasses the selection of representative partial differential equations, detailed experimental configurations, and the adopted evaluation metrics. A comprehensive comparative analysis will be conducted across four models: the standard PINN, DBAW-PINN, PIKAN, and the proposed DBAW-PIKAN. Our evaluation primarily focuses on the discrepancies in their training accuracy, convergence rates, and computational efficiency.
@@ 3.1 Klein-Gordon Equation
The Klein-Gordon equation is a fundamental equation in relativistic quantum mechanics and quantum field theory. As a nonlinear partial differential equation, it finds extensive applications in fields such as particle physics and nonlinear optics. The Klein-Gordon equation is formulated as:
@@
We compare the performance of four models: PINN, DBAW-PINN, PIKAN, and DBAW-PIKAN. For the MLP-based architectures (PINN and DBAW-PINN), the network consists of 7 hidden layers with 64 neurons each. For the KAN-based architectures (PIKAN and DBAW-PIKAN), the structure is configured with 4 hidden layers, each with a width of 20, a B-spline order of 4, and a grid size of 20. All models are trained using the Adam optimizer for 50,000 iterations. Table 1 provides a detailed performance comparison of the four methods for solving the Klein-Gordon equation. Quantitatively, the DBAW-PIKAN method achieves the lowest relative $L_{2}$ error of $8.88\times 10^{-4}$. In contrast, the standard PINN model yields the highest error at $3.05\times 10^{-2}$. The introduction of the Dynamic Balancing Adaptive Weighting (DBAW) strategy reduces the error of DBAW-PINN to $1.17\times 10^{-2}$, validating the effectiveness of the self-adaptive weighting mechanism. Furthermore, when the architecture is replaced by KAN, the PIKAN model further reduces the error to $7.73\times 10^{-3}$, showcasing the superior function approximation capabilities of KAN. Ultimately, the DBAW-PIKAN model, which integrates both improvements, achieves a precision nearly an order of magnitude higher than that of PIKAN. This fully demonstrates the synergistic effect between the PIKAN architecture and the DBAW strategy.
@@ 3.2 Burgers
$u(x,0)$ $=g(x),\quad x\in\Omega$
$u(-1,t)$ $=h_{1}(t),\quad t\in[0,T]$
$u(1,t)$ $=h_{2}(t),\quad t\in[0,T]$
While analytical solutions for this equation are generally unavailable, numerical approximations can be obtained using various computational methods [raissi2019physics].
Consistent with the previous experimental configurations, we evaluate four models: PINN, DBAW-PINN, PIKAN, and DBAW-PIKAN. Specifically, the MLP-based models (PINN and DBAW-PINN) utilize an architecture of [2, 64, 64, 64, 64, 64, 64, 1], while the KAN-based models (PIKAN and DBAW-PIKAN) employ a [2, 20, 20, 20, 1] structure with a B-spline order of 4 and a grid size of 20. All models are trained for 50,000 iterations using the Adam optimizer. Table 2 summarizes the quantitative results for the Burgers equation. The standard PINN achieves a relative $L_{2}$ error of $1.23\times 10^{-3}$. Notably, DBAW-PINN exhibits suboptimal performance on this problem, with the error escalating to $2.71\times 10^{-1}$. This suggests that for certain nonlinear problems, adaptive weighting strategies may interfere with the training process if not properly compatible with the specific MLP architecture. In contrast, the PIKAN model demonstrates superior performance, reducing the error to $1.16\times 10^{-3}$, which further underscores the potential of KAN in fitting complex nonlinear functions. Ultimately, the proposed DBAW-PIKAN method yields the highest precision with an error of $\mathbf{4.70\times 10^{-4}}$. This significant improvement over the standard PIKAN validates the effectiveness of integrating the dynamic adaptive weighting strategy with the KAN architecture. The numerical solutions are visually compared in Fig. 6: (a) The PINN prediction shows visible deviations from the reference solution near the shock wave region. (b) The DBAW-PINN prediction severely deviates from the reference. (c) The PIKAN prediction is in good overall agreement with the reference. (d) The DBAW-PIKAN solution is visually the most accurate, particularly in precisely capturing the shock front, with its corresponding error map exhibiting the lowest overall error level. To further elucidate the dynamic optimization mechanism of DBAW-PIKAN when solving the nonlinear Burgers equation, Fig. 5 illustrates the evolution trajectories of key parameters during training. As shown in Fig. 5(a), the boundary condition weight $\lambda_{bc}$ (orange curve) escalates rapidly during the initial phase. This aligns with the characteristic training behavior of PINNs, where simple functional relationships on the boundaries are typically easier for the network to fit than the complex interior PDE residuals. Without intervention, such a rapid increase in weight would cause the optimizer to over-focus on the boundary loss, thereby neglecting the learning of the underlying PDE physics.
@@
(b) Evolution of the dynamic upper bound
@@
Figure 5: Parameter evolution during the training process for the Burgers equation: (a) trajectories of adaptive weights ($\lambda_{r}$, $\lambda_{ic}$, $\lambda_{bc}$) over training epochs; (b) evolution of the dynamic upper bound parameter $\gamma(t)$ over training epochs; and (c) evolution of individual loss components over training epochs.

### v2

@@ 2.3 Dynamic Adaptive Loss Weighting Function
While PINNs have proven effective for many simple problems, they often face significant challenges when applied to more complex ones. A primary contributing factor is that traditional PINNs use fixed weights to balance these loss terms, but their performance is highly dependent on the choice of weights and requires tedious manual tuning. Therefore, selecting appropriate loss weights to balance the various components of the loss function can effectively accelerate the convergence speed and improve the accuracy of PINNs[16][31]. Inspired by the uncertainty weighting method in multi-task learning[13], we can treat each loss term—which typically includes the PDEs residual, Boundary Conditions, and Initial Conditions—as a separate task. Thus, the process of solving a PDE can be viewed as a multi-task learning problem. By using homoscedastic uncertainty as the basis for loss weighting in a multi-task learning problem, we note that homoscedastic uncertainty is an impromptu uncertainty that does not depend on the input data. It is not a model output but a quantity that remains constant across all input data and varies between different tasks. Therefore, it can be described as task-dependent uncertainty. The implementation process is as follows. We assume that the PDE solution follows a Gaussian distribution, and its likelihood function can be expressed as:
$p(u|\hat{u},\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\|u-\hat{u}\|^{2}}{2\sigma^{2}}\right).$ (14)
@@
$p(u|\hat{u},\sigma_{r}^{2},\sigma_{b}^{2},\sigma_{i}^{2})=\prod_{j\in\{r,b,i\}}\frac{1}{\sqrt{2\pi\sigma_{j}^{2}}}\exp\left(-\frac{\mathcal{L}_{j}}{2\sigma_{j}^{2}}\right),$ (15)
@@
$\begin{split}&-\log p(u|\hat{u},\sigma_{r}^{2},\sigma_{b}^{2},\sigma_{i}^{2})\\ &=-\sum_{j\in\{r,b,i\}}\log\biggl[\frac{1}{\sqrt{2\pi\sigma_{j}^{2}}}\exp\biggl(-\frac{\|u-\hat{u}\|^{2}}{2\sigma_{j}^{2}}\biggr)\biggr]\\ &=\sum_{j\in\{r,b,i\}}\biggl(\frac{1}{2\sigma_{j}^{2}}\mathcal{L}_{j}+\frac{1}{2}\log\sigma_{j}^{2}+\frac{1}{2}\log(2\pi)\biggr).\end{split}$ (16)
@@
$\mathcal{L}_{\text{base}}(W,\sigma_{r}^{2},\sigma_{b}^{2},\sigma_{i}^{2})=\sum_{j\in\{r,b,i\}}\left(\frac{1}{2\sigma_{j}^{2}}\mathcal{L}_{j}(W)+\frac{1}{2}\log\sigma_{j}^{2}\right).$ (17)
The first term in Eq.(17) represents the weighted loss function for each task, while the second term is a regularization term that penalizes tasks with high uncertainty. Minimizing $\mathcal{L}_{\text{base}}$ in Eq.(17) is equivalent to maximizing the joint probability distribution of all losses, which facilitates the adaptive allocation of appropriate weight coefficients to the various components of the loss, ensuring a balance among them. The parameter $\sigma_{j}^{2}$ represents the uncertainty associated with each task. If a task is easy to learn and exhibits low uncertainty, it will be assigned a higher weight; conversely, a task with high uncertainty will receive a lower weight. In summary, the uncertainty-based loss function is:
@@
$\quad+\frac{1}{2}\log\sigma_{r}^{2}+\frac{1}{2}\log\sigma_{bc}^{2}+\frac{1}{2}\log\sigma_{ic}^{2}.$
However, in multi-task PDE solving, different loss terms (PDE residual, boundary conditions, initial conditions) often have different orders of magnitude and convergence characteristics. In the early stages of training, certain loss terms may dominate the gradient updates. If the gradient magnitude of some loss terms is too large, it can lead to training instability and weight imbalance, causing the weights of some tasks to grow excessively, thereby suppressing the learning of other tasks and causing the model to prematurely fall into local optima for certain tasks. Therefore, inspired by [17], this paper introduces a dynamic upper bound weight constraint method. Its core mathematical expression is:
$\gamma(t)=\gamma_{\text{max}}\cdot\exp(-\alpha t)+\gamma_{\text{min}},$ (19)
@@
$\lambda_{j}=\frac{1}{\sigma_{j}^{2}+\frac{1}{\gamma(t)}+\epsilon},$ (20)
where $\epsilon=10e^{-12}$ is a small constant added for numerical stability. When $\gamma(t)$ is large, $\frac{1}{\gamma(t)}\rightarrow 0$, and the weight is mainly controlled by $\sigma_{j}^{2}$; when $\gamma(t)$ is small, the term $\frac{1}{\gamma(t)}$ becomes significant, constraining the range of weight changes.
Finally, we obtain the extended constrained adaptive loss function:
$\mathcal{L}_{\text{adaptive}}=\sum_{j\in\{r,b,i\}}\left(\lambda_{j}\mathcal{L}_{j}+\log\left(\sigma_{j}^{2}+\frac{1}{\gamma(t)}\right)\right).$ (21)
@@ 2.4 DBAW-PIKAN
To simultaneously address the limitations in network expressivity and the imbalance in multi-objective optimization within the PINNs framework, we propose an integrated DBAW-PIKAN framework. This framework creatively integrates the high-expressivity PIKAN architecture with the DBAW self-adaptive weighting strategy.
This integration is not a mere superposition of two independent improvements but is fundamentally rooted in the synergy between them.On one hand, the PIKAN architecture, as constructed in Section 2.2, leverages the superior function approximation capabilities of Kolmogorov-Arnold Networks to provide the potential for accurate PDE solutions, effectively mitigating the spectral bias inherent in traditional MLPs [30]. However, the potential of a high-expressivity architecture cannot be fully realized if the training process is unstable. As an implementation of PINNs, PIKAN still inherits the core challenge of imbalanced training dynamics caused by multi-objective loss functions.On the other hand, the DBAW strategy proposed in Section 2.3 provides stability to the optimization process by dynamically balancing the contributions of various loss components. This effectively prevents certain loss terms from dominating the optimization and suppressing the learning of other physical constraints. However, experimental results (see Section 3.2) indicate that the DBAW strategy is not a universal ”plug-and-play” module. When DBAW is combined with a traditional MLP architecture (i.e., the DBAW-PINN model), it exhibits suboptimal performance in solving the nonlinear Burgers equation. This suggests that the efficacy of the DBAW strategy is highly dependent on a compatible and sufficiently expressive underlying network architecture.
@@
Input: $\gamma_{\max},\gamma_{\min},\alpha,N_{int},N_{bc},T$
@@
Step1: Initialize the Kolmogorov-Arnold Network (KAN) model $f_{\theta}$ and adaptive weight parameters $\log\sigma_{r}\leftarrow 0,\log\sigma_{u}\leftarrow 0$.
Step2: Initialize the optimizers $\text{Adam}_{\theta}$ and $\text{Adam}_{\sigma}$ with designated learning rates for model parameters and weights respectively.
Step3: Execute the training loop for $t=1,\dots,T$:
  1. (a)
    Sample interior points $\{(x_{i},y_{i})\}_{i=1}^{N_{int}}$ and boundary points $\{(x_{j}^{b},y_{j}^{b})\}_{j=1}^{N_{bc}}$.
  2. (b)
    Compute the dynamic upper bound $\gamma(t)=\gamma_{\max}\cdot\exp(-\alpha t)+\gamma_{\min}$ and evaluate PDE residual loss $\mathcal{L}_{PDE}$ and boundary loss $\mathcal{L}_{BC}$.
  3. (c)
    Calculate the adaptive weights $\lambda_{r}$ and $\lambda_{u}$ as:
    $\lambda_{r}=\min\left\{\frac{1}{\sigma_{r}^{2}+\frac{1}{\gamma(t)}},\gamma(t)\right\},\quad\lambda_{u}=\min\left\{\frac{1}{\sigma_{u}^{2}+\frac{1}{\gamma(t)}},\gamma(t)\right\}.$
  4. (d)
    Update the KAN model parameters $\theta$ and weight parameters $\log\sigma$:
    $\theta_{t+1}$ $=\text{Adam}_{\theta}(\mathcal{L}_{Total},\theta_{t})$
    $\log\sigma_{r,t+1},\log\sigma_{u,t+1}$ $=\text{Adam}_{\sigma}(\mathcal{L}_{Total},\log\sigma_{r,t},\log\sigma_{u,t})$
Return: The optimized model parameters $\theta^{*}$ and the final adaptive weights $\log\sigma_{r}^{*},\log\sigma_{u}^{*}$.
@@ 3 Experiments
To quantify the prediction error of the models, we employ the relative $L_{2}$ error as the performance metric, which is defined as:
$L_{2}=\frac{\sqrt{\sum_{i=1}^{N}\left|\hat{u}\left(x_{i},t_{i}\right)-u\left(x_{i},t_{i}\right)\right|^{2}}}{\sqrt{\sum_{i=1}^{N}\left|u\left(x_{i},t_{i}\right)\right|^{2}}},$ (22)
@@
The objective of this section is to evaluate the performance of the proposed DBAW self-adaptive loss strategy and the enhanced PIKAN framework through a series of numerical experiments. The experimental design encompasses the selection of representative partial differential equations, detailed experimental configurations, and the adopted evaluation metrics. A comprehensive comparative analysis will be conducted across four models: the standard PINNs, DBAW-PINN, PIKAN, and the proposed DBAW-PIKAN. Our evaluation primarily focuses on the discrepancies in their training accuracy, convergence rates, and computational efficiency.
@@ 3.1 Klein-Gordon Equation
The Klein-Gordon equation describes the dynamics of spinless particles within the frameworks of relativistic quantum mechanics and quantum field theory. As a nonlinear partial differential equation, it finds extensive applications in fields such as particle physics and nonlinear optics. The Klein-Gordon equation is formulated as:
@@
We compare the performance of four models: PINN, DBAW-PINN, PIKAN, and DBAW-PIKAN. For the MLP-based architectures (PINN and DBAW-PINN), the network consists of 7 hidden layers with 64 neurons each. For the KAN-based architectures (PIKAN and DBAW-PIKAN), the structure is configured with 4 hidden layers, each with a width of 20, a B-spline order of 4, and a grid size of 20. All models are trained using the Adam optimizer for 50,000 iterations. Table 1 provides a detailed performance comparison of the four methods for solving the Klein-Gordon equation. Quantitatively, the DBAW-PIKAN method achieves the lowest relative $L_{2}$ error of $8.88\times 10^{-4}$. In contrast, PINN yields the highest error at $3.05\times 10^{-2}$. The introduction of the Dynamic Balancing Adaptive Weighting (DBAW) strategy reduces the error of DBAW-PINN to $1.17\times 10^{-2}$, validating the effectiveness of the self-adaptive weighting mechanism. Furthermore, when the architecture is replaced by KAN, the PIKAN model further reduces the error to $7.73\times 10^{-3}$, showcasing the superior function approximation capabilities of KAN. Ultimately, the DBAW-PIKAN model, which integrates both improvements, achieves a precision nearly an order of magnitude higher than that of PIKAN. This fully demonstrates the synergistic effect between the PIKAN architecture and the DBAW strategy.
@@ 3.2 Burgers
$u(x,0)$ $=g(x),\quad x\in\Omega,$
$u(-1,t)$ $=h_{1}(t),\quad t\in[0,T],$
$u(1,t)$ $=h_{2}(t),\quad t\in[0,T].$
While analytical solutions for this equation are generally unavailable, numerical approximations can be obtained using various computational methods [20].
Consistent with the previous experimental configurations, we evaluate four models: PINN, DBAW-PINN, PIKAN, and DBAW-PIKAN. Specifically, the MLP-based models (PINN and DBAW-PINN) utilize an architecture of [2, 64, 64, 64, 64, 64, 64, 1], while the KAN-based models (PIKAN and DBAW-PIKAN) employ a [2, 20, 20, 20, 1] structure with a B-spline order of 4 and a grid size of 20. All models are trained for 50,000 iterations using the Adam optimizer. Table 2 summarizes the quantitative results for the Burgers equation. The standard PINN achieves a relative $L_{2}$ error of $1.23\times 10^{-3}$. Notably, DBAW-PINN exhibits suboptimal performance on this problem, with the error escalating to $2.71\times 10^{-1}$. This suggests that for certain nonlinear problems, adaptive weighting strategies may interfere with the training process if not properly compatible with the specific MLP architecture.
@@
In contrast, the PIKAN model demonstrates superior performance, reducing the error to $1.16\times 10^{-3}$, which further underscores the potential of KAN in fitting complex nonlinear functions. Ultimately, the proposed DBAW-PIKAN method yields the highest precision with an error of $\mathbf{4.70\times 10^{-4}}$. This significant improvement over the standard PIKAN validates the effectiveness of integrating the dynamic adaptive weighting strategy with the KAN architecture. The numerical solutions are visually compared in Fig. 6: (a) The PINN prediction shows visible deviations from the reference solution near the shock wave region. (b) The DBAW-PINN prediction severely deviates from the reference. (c) The PIKAN prediction is in good overall agreement with the reference. (d) The DBAW-PIKAN solution is visually the most accurate, particularly in precisely capturing the shock front, with its corresponding error map exhibiting the lowest overall error level. To further elucidate the dynamic optimization mechanism of DBAW-PIKAN when solving the nonlinear Burgers equation, Fig. 5 illustrates the evolution trajectories of key parameters during training. As shown in Fig. 5(a), the boundary condition weight $\lambda_{bc}$ (orange curve) escalates rapidly during the initial phase. This aligns with the characteristic training behavior of PINNs, where simple functional relationships on the boundaries are typically easier for the network to fit than the complex interior PDE residuals. Without intervention, such a rapid increase in weight would cause the optimizer to over-focus on the boundary loss, thereby neglecting the learning of the underlying PDE physics.
@@
(b) Evolution of the dynamic upper bound $\gamma(t)$
@@
Figure 5: Parameter evolution during the training process for the Burgers equation.